A fire fighters crew uses water cannon at 25m/s at a fixed angle of 53.0 degree above horizontal. The fire fighters want to direct the water to a blaze that is 10.0m above ground level. How far from the blaze should they position their cannon?


1
Expert's answer
2021-06-07T09:35:22-0400

Explanations & Calculations


  • Careful inspection of the situation along with the motion equation "\small (x-x_0 )=v_0t+\frac{1}{2}at^2" express that there exist two sets of values for the time that correspond to a given displacement.
  • Then, applying this to the given situation yields,

"\qquad\qquad
\begin{aligned}
\small \uparrow10m&=\small 25\sin53 t+\frac{1}{2}\times(-9.8ms^{-2})t^2\\
\small t&=\begin{cases}
\small 3.489\,s\\
\small 0.585\,s
\end{cases}
\end{aligned}"


  • It should be the larger value that fits the given situation as the trajectory of the water column is a projectile.



  • Then we can calculate from how far the water jet should be aimed.

"\qquad\qquad
\begin{aligned}
\small \to (x-x_0)&=\small v_0 t\\
\small x&=\small 25\cos 53\times 3.489\\
&=\small \bold{52.493\,m}
\end{aligned}"